# New Tight Bounds for SGD without Variance Assumption: A Computer-Aided Lyapunov Analysis

### v1

@@ E.2 Converting PEP solutions into mathematical proofs
If we have a primal-dual optimal solution $(\overline{G}_{t},\overline{F}_{t},\overline{\Lambda}_{t},(\overline{\lambda}_{t,*}^{(i)}),(\overline{\lambda}_{*,t}^{(i)}))$, it holds that, for all $G_{t}\in\mathcal{S}^{2m}$ and $F_{t}\in\mathbb{R}^{2m}$,
$L(\overline{G}_{t},\overline{F}_{t},\overline{\Lambda}_{t},(\overline{\lambda}_{t,*}^{(i)}),(\overline{\lambda}_{*,t}^{(i)}))\leq L(G_{t},F_{t},\overline{\Lambda}_{t},(\overline{\lambda}_{t,*}^{(i)}),(\overline{\lambda}_{*,t}^{(i)})).$
@@
$L(\overline{G}_{t},\overline{F}_{t},\overline{\Lambda}_{t},(\overline{\lambda}_{t,*}^{(i)}),(\overline{\lambda}_{*,t}^{(i)}))=-\operatorname{Tr}(\Delta_{t}\overline{G}_{t})-\overline{F}_{t}\tilde{\Delta}=B_{t}.$
By strong duality (see Section E.3), we know that the existence of a primal-dual solution implies that $B_{t}=0$. As such, it holds that, for all $G_{t}\in\mathcal{S}^{2m}$ and $F_{t}\in\mathbb{R}^{2m}$,
$0\leq L(G_{t},F_{t},\overline{\Lambda}_{t},(\overline{\lambda}_{t,*}^{(i)}),(\overline{\lambda}_{*,t}^{(i)})),$
@@
$\operatorname{Tr}(\Delta_{t}G_{t})+F_{t}\tilde{\Delta}+\operatorname{Tr}(\overline{\Lambda}_{t}G_{t})$ $\leq\sum_{i=1}^{m}\overline{\lambda}_{t,*}^{(i)}\left(F_{t}\bm{f}_{m+i}-F_{t}\bm{f}_{i}+\operatorname{Tr}(A_{t,*}^{(i)}G_{t})\right)$ (11)
$+\sum_{i=1}^{m}\overline{\lambda}_{*,t}^{(i)}\left(F_{t}\bm{f}_{i}-F_{t}\bm{f}_{m+i}+\operatorname{Tr}(A_{*,t}^{(i)}G_{t})\right).$
As $(\overline{G}_{t},\overline{F}_{t},\overline{\Lambda}_{t},(\overline{\lambda}_{t,*}^{(i)}),(\overline{\lambda}_{*,t}^{(i)}))$ is primal-dual optimal, in particular it is primal-dual feasible, and hence $\overline{\Lambda}_{t}\succeq 0$ and $\overline{\lambda}_{t,*}^{(i)},\overline{\lambda}_{*,t}^{(i)}\geq 0$ for all $i=1,\ldots,m$. In particular, if $G_{t}$ and $F_{t}$ are generated by functions $f_{i}\in\mathcal{F}_{\mu,L}$ through Equations (10), it holds that the right-hand side of Equation (11) is nonpositive by Theorem E.1. By recalling that $E_{t+1}-E_{t}=\operatorname{Tr}(\Delta_{t}G_{t})+F_{t}\tilde{\Delta}$,
$E_{t+1}-E_{t}\leq-\operatorname{Tr}(\overline{\Lambda}_{t}G_{t})\leq 0,$
where the latter follows since $\overline{\Lambda}_{t},G_{t}\succeq 0$. Specifically, one can prove that $E_{t+1}-E_{t}\leq 0$ by rewriting it in the form of (11), using the obtained dual optimal solution, and writing $\operatorname{Tr}(\overline{\Lambda}_{t}G_{t})$ as an appropriate sum of squares based on the Cholesky factorization of $\overline{\Lambda}_{t}$.
E.3 Strong duality
Theorem E.2.
@@ E.3 Strong duality
$+\left(\tilde{\Delta}+\sum_{i=1}^{m}\lambda_{*,t}^{i}(\bm{f}_{m+i}-\bm{f}_{i})+\sum_{i=1}^{m}\lambda_{*,t}^{i}(\bm{f}_{i}-\bm{f}_{m+i})\right)F_{t},$
from which we readily recover Problem (Dual) as the dual. In order to show that strong duality holds, we employ a standard Slater argument (boyd_convex_2023) and construct a feasible point $(G_{t},F_{t})$ to Problem (Primal) such that $G_{t}\succ 0$.
To do so, we follow the same ideas as in (taylor_smooth_2017, Theorem 6). Specifically, leveraging the discussion in Section E.1, we shall construct $m$ functions $f_{i}\in\mathcal{F}_{\mu,L}$ and points $x_{t},x_{*}\in\mathbb{R}^{d}$ such that the matrix $P_{t}$ given by Equation (10a) is upper triangular with positive entries on its diagonal. As such, assuming $d\geq 2m$, it will hold that $G_{t}=P_{t}^{T}P_{t}\succ 0$. We assume without loss of generality that $x_{*}=0$, and set $d=2m$.
@@
where $Q_{i}^{1}\in\mathbb{R}^{(m-1)\times(m-1)}$ is given by
$Q_{i}^{1}=\text{diag}(2)+3\bm{e}_{i}\bm{e}_{i}^{T},$
and $Q\in\mathbb{R}^{(m+1)\times(m+1)}$ is a tridiagonal matrix whose diagonal elements are $4$ and whose off diagonal elements are $-1$. We note that $\sigma(Q_{i}^{1})\subset[2,5]$, and that, by the Greshgorian Theorem (quarteroni_numerical_2007), $\sigma(Q-\bm{e}_{i+1}\bm{e}_{i+1}^{T})\subset(1,6)$. Specifically, $\sigma(Q_{i})\subset(1,6)$ for all $i=1,\ldots,m-1$.
We define $x_{t}=[x^{1},x^{2}]$ with $x^{1}\in\mathbb{R}^{m-1}$ and $x^{2}\in\mathbb{R}^{m+1}$. Then
$\nabla f_{i}(x_{t})=g_{t}^{(i)}=2\sum_{j=1,j\neq i}^{m-1}\bm{e}_{j}x^{1}_{j}+5\bm{e}_{i}x^{1}_{i}+Qx^{2}-\bm{e}_{m+i}x^{2}_{i+1}+b_{i}.$
@@
$b_{i}=-2\sum_{j=1,j\neq i}^{m-1}\bm{e}_{j}x^{1}_{j}+\bm{e}_{m+i}x^{2}_{i+1}-\bm{e}_{m},$
and we impose that $x^{2}>0$ and that $Qx^{2}=[1,0,\ldots,0]\in\mathbb{R}^{m+1}$ (may be shown possible by a simple induction argument). This thus gives us that
$g_{t}^{(i)}=5\bm{e}_{i}x^{1}_{i}.$
Moreover, as $x_{*}=0$, we have that $g_{*}^{(i)}=b_{i}$. We additionally impose $x^{1}>0$. Note that $g_{t}^{(i)}$ has a positive entry in its $i$th component and 00s everywhere else. Moreover, $g_{*}^{(i)}$ has a positive component in its $m+i$th component and 00 in all subsequent components. The vector $x_{t}-x_{*}$ also has a positive entry in component $2m$. In order to show that $P_{t}$ is upper triangular with positive diagonal entries, we thus only need to construct $g_{t}^{(m)}$.
@@
$\tilde{f}_{i}(x)=(f_{i}(x)-\tfrac{1}{2}\|x\|^{2})\frac{L-\mu}{5}+\mu\tfrac{1}{2}\|x\|^{2}.$
@@
As $f_{i}\in\mathcal{F}_{1,6}$, we have $\tilde{f}_{i}\in\mathcal{F}_{\mu,L}$. Moreover, this transformation preserves the property on $P_{t}$, hence constructing a Slater point for all $(\mu,L)$. ∎
Appendix F More details on the numerical experiments
This section complements the numerics included in Section 4. Specifically, we compare the theoretical bounds obtained in Sections C and D with the numerical bounds obtained through the Performance Estimation Problem (PEP) methodology, as explained in Section E.
All our experiments were run with $L=1$, which can always be assumed to hold, up to rescaling the functions defining our problem.
F.1 Convex case, non-optimal step-size
Bias term
Figure 5 showcases the tightness of our theoretical bound on the bias term in the convex setting with non-optimal step-size by comparing it to the worst-case performance computed using the PEP methodology. Across iteration horizons $T=2,20,100$, the theoretical bound closely matches the numerical results obtained via PEP, indicating that the bound is not only valid but also tight. This agreement confirms that our theoretical analysis accurately captures the conditional worst-case behavior of SGD within the range of analyses permitted by our Lyapunov energies, and that the bound cannot be significantly improved under the given assumptions.
Figure 5: Theoretical and numerical bias term in the convex setting for the full range of step-sizes.
Variance term under tight bias
As a bi-product, we are able to exhibit upper bounds on the variance term, once the optimal bias has been established. In the short step-size range, depicted in Figure 6, we observe that our theoretical bounds closely align with those obtained through the PEP methodology, strongly suggesting that the bound cannot be further improved within the constraints of our Lyapunov analysis. In the large step-size range, shown in Figure 7, we observe similar results, although our results widely suffer from numerical instabilities.
Figure 6: Theoretical and numerical average variance in the convex setting for short step-sizes.
Figure 7: Theoretical and numerical average variance in the convex setting for large step-sizes.
F.2 Convex case, optimal step-size
We note that the results in the figure below are not tight due to our restriction that $a_{t}\equiv 1$, which is not the case for $\varepsilon>0$. Nevertheless, as $T\to\infty$, the PEP bound seems to converge to our (asymptotic) bound. The plot is generated by fixing a rate of $\rho=2\gamma-\varepsilon$, and minimizing the average variance.
Figure 8: Numerical and theoretical variance for optimal step-size and sub-optimal bias term.
Combining Figures 6 and 8, we note that it seems impossible to obtain a finite variance for the optimal bias term at the optimal step-size $\gamma L=1$. This confirms our discussion in Section C.2.
F.3 Strongly convex case, non-optimal step-size
Bias term
As in the convex setting, we hint towards the tightness of our theoretical bound on the bias in the strongly convex setting for non-optimal step-sizes in Figure 9. The overlap of the theoretical curve and the numerically generated one once again confirms our tightness.
Figure 9: Theoretical and numerical bias term in the strongly convex setting with $\mu=0.25$.
Variance term under tight bias
We moreover assess the tightness of the variance term, by first fixing the bias term to be the observed optimal, and then minimizing the total variance. The results are shown in Figure 10, where the numerics confirm the theoretical claims that we cannot obtain a finite variance term for the optimal bias term in the optimal step-size.
Figure 10: Theoretical and numerical variance in the strongly convex setting with $\mu=0.25$.
We define the relative error $\tau$ between the theoretical and numerical average variance as
$\tau=\frac{\overline{e}_{\text{opt}}-\overline{e}_{\text{theory}}}{\overline{e}_{\text{theory}}},$
where $\overline{e}_{\text{theory}}$ is the theoretical average variance term obtained in Theorem 3.3. Figure 11 shows this relative error for various values of $T$ over a range of values of step-sizes $\gamma L$ and strong convexity parameters $\mu$.
Figure 11: Relative errors $\tau$ in optimal average variance
F.4 Strongly convex case, optimal step-size
In Figure 12 we plot the numerical and theoretical variance term upon fixing an $\varepsilon$-tight bias term. Our bounds might not appear tight, but we suspect this to be due to numerical instabilities. To emphasize this, we include the same results using a different solver in Figure 13, whose results are widely different. We leave this for further investigation.
Figure 12: Theoretical and numerical variance in the strongly convex setting with $\mu=0.25$, in the optimal step-size (using MOSEK).
Figure 13: Theoretical and numerical variance in the strongly convex setting with $\mu=0.25$, in the optimal step-size (using CLARABEL).
F.5 Details on numerical instability
We note that the problems solved numerically are high-dimensional and grow with the number of iterations $T$. As such, for $T$ large, we suffer from numerical instability. Figure 14 shows that a slight change in tolerance provided to the solver might drastically modify the obtained solution. We shall not delve further into this.
Figure 14: Example of numerical instability for $(L,\mu)=(1,0.5)$.
Appendix G SGD with non-uniform sampling and mini-batching
@@
All the main results in our paper can be applied to variants of (SGD), such as mini-batching or nonuniform sampling. This is because these variants can be seen as instances of (SGD) applied to different but equivalent problems. Our results can be applied as a black-box, and the results follow by computing the main constants of the equivalent problem, namely $L,\mu,\sigma_{*}^{2}$. This strategy was described in gower_sgd_2019, which contains all the formulas needed to compute those constants. For the sake of completeness, we include some details for the two aforementioned instances below.
G.1 SGD with non-uniform sampling
@@
where the expectation is taken with respect $\mathcal{P}$, the distribution whose density is given by the probabilities $p_{i}$. Note that the function $f$ remains unchanged, meaning that the set of minimizers is the same. We can then compute (gower_sgd_2019, Propositions 3.7 & 3.8)) the following quantities:
  • $L$, which is the maximum among the constants $\operatorname{Lip}(\nabla\hat{f}_{i})$, so that we can say that each $\hat{f}_{i}$ is $L$-smooth. Clearly,
    $L=\max\limits_{i=1,\dots,n}\ \frac{L_{i}}{np_{i}}.$
  • $\mu$, which, for the same reasons as above, can be computed as
    $\mu=\min\limits_{i=1,\dots,n}\ \frac{\mu_{i}}{np_{i}}.$
  • $\sigma_{*}^{2}$, which is given by
    $\mathbb{E}_{{\mathcal{P}}}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}=\sum_{i=1}^{n}p_{i}\|\nabla\hat{f}_{i}(x_{*})\|^{2}=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{np_{i}}\|\nabla\hat{f}_{i}(x_{*})\|^{2}.$
If we now impose that $\gamma L\in(0,2)$, we can apply all our results and obtain bounds on the iterates of (SGDp) depending on the above constants.
G.2 SGD with mini-batching
@@
To be able to apply our results, we need to verify that this new problem verifies our assumptions. More precisely, all we need is for it to verify (SEC) for some $L$ and some $\mu$. This can be calculated using the results in gower_sgd_2019.
Lemma G.1.
Consider Problem A.1 with $\mathcal{I}$ finite. Assume that each function $f_{i}$ is $\mu_{i}$-convex and $L_{i}$-smooth. Then the family of functions $(f_{B})_{B\in\text{batch}_{b}}$ verifies (SEC) with constants
  • $L=(1-r)\operatorname{Lip}(\nabla f)+r\max_{i}L_{i}$, with $r=\tfrac{n(b-1)}{b(n-1)}\in[0,1]$,
  • $\mu=\min_{i}\mu_{i}$.
Moreover, we have that
@@
$\sigma_{*}^{2}=\mathbb{E}_{{\mathcal{B}}}\mathopen{}\left[{\|\nabla f_{B}(x_{*})\|^{2}}\right]\mathclose{}=\frac{n-b}{nb(n-1)}\sum_{i=1}^{n}\|\nabla f_{i}(x_{*})\|^{2}.$
Proof.
Let $\mu\coloneqq\min_{i}\mu_{i}$, such that each $f_{i}$ is $\mu$-convex. Similarly, because each $\hat{f}_{B}$ is an average of $f_{i}$’s, we know that each $\hat{f}_{B}$ is $\mu$-convex. We can therefore define
$g=f-\tfrac{\mu}{2}\|\cdot\|^{2},\quad g_{i}=f_{i}-\tfrac{\mu}{2}\|\cdot\|^{2}\quad\text{ and }\quad\hat{g}_{B}=\hat{f}_{B}-\tfrac{\mu}{2}\|\cdot\|^{2}$
all of which are convex functions. By definition, we know that each $\hat{g}_{B}$ is an average of functions $g_{i}$ which are each $(L_{i}-\mu)$-smooth and convex. Invoking (gower_sgd_2019, Section G) shows that the family $(\hat{g}_{B})_{B\in\text{batch}_{b}}$ verifies (EC) for constants $(\mathcal{L},0)$ where
$\mathcal{L}=(1-r)\operatorname{Lip}(\nabla g)+r\max_{i}\operatorname{Lip}(\nabla g_{i}),\text{ with }r=\tfrac{n(b-1)}{b(n-1)}\in[0,1].$
Using Lemma G.2 which can be found below, we see that this is equivalent to the family $(\hat{f}_{B})_{B\in\text{batch}_{b}}$ verifying (EC) for constants $(\mathcal{L}+\mu,\mu)$. Because $\operatorname{Lip}(\nabla g)=\operatorname{Lip}(\nabla f)-\mu$ and $\operatorname{Lip}(\nabla g_{i})=\operatorname{Lip}(\nabla f_{i})-\mu$, we conclude by setting $L=\mathcal{L}+\mu$. ∎
Lemma G.2.
Consider Problem A.1. Let $L\geq\mu\geq 0$, and let $(x,y)\in\mathcal{H}$ be fixed. Then the following are equivalent:
  1. 1.
    the family of functions $(f_{i})_{i\in\mathcal{I}}$ verify (EC) with constants $(L,\mu)$;
  2. 2.
    the family of functions $(g_{i})_{i\in\mathcal{I}}$ verify (EC) with constants $(L-\mu,0)$, with $g_{i}=f_{i}-\tfrac{\mu}{2}\|\cdot\|^{2}$.
Proof.
It suffices to show that, if $\hat{f},\hat{g}\colon\mathbb{R}^{d}\to\mathbb{R}$ are such that $\hat{g}=\hat{f}-\tfrac{\mu}{2}\|\cdot\|^{2}$, then
$\tfrac{1}{2}\|\nabla\hat{f}(y)$ $-\nabla\hat{f}(x)\|^{2}+\tfrac{\mu L}{2}\|y-x\|^{2}-(L-\mu)\big{[}\hat{f}(y)-\hat{f}(x)\big{]}-\langle\mu\nabla\hat{f}(y)-L\nabla\hat{f}(x),y-x\rangle$
$=\tfrac{1}{2}\|\nabla\hat{g}(y)-\nabla\hat{g}(x)\|^{2}-(L-\mu)\big{[}\hat{g}(y)-\hat{g}(x)\big{]}+(L-\mu)\langle\nabla\hat{g}(x),y-x\rangle.$ (12)
We start from the right-hand side, substituting $\hat{g}(z)=\hat{f}(z)-\tfrac{\mu}{2}\|z\|^{2}$ and $\nabla\hat{g}(z)=\nabla\hat{f}(z)-\mu z$ for $z=x$ and $z=y$, to obtain
$\tfrac{1}{2}\|\nabla\hat{g}(y)$ $-\nabla\hat{g}(x)\|^{2}-(L-\mu)\big{[}\hat{g}(y)-\hat{g}(x)\big{]}+(L-\mu)\langle\nabla\hat{g}(x),y-x\rangle$
=\ = $\tfrac{1}{2}\|\nabla\hat{f}(y)-\nabla\hat{f}(x)\|^{2}-\mu\langle\nabla\hat{f}(y)-\nabla\hat{f}(x),y-x\rangle+\tfrac{\mu^{2}}{2}\|y-x\|^{2}$
$-(L-\mu)\big{[}\hat{f}(y)-\hat{f}(x)\big{]}+\tfrac{\mu}{2}(L-\mu)\big{[}\|y\|^{2}-\|x\|^{2}\big{]}$
$+(L-\mu)\langle\nabla\hat{f}(x),y-x\rangle-\mu(L-\mu)\langle x,y-x\rangle$
=\ = $\tfrac{1}{2}\|\nabla\hat{f}(y)-\nabla\hat{f}(x)\|^{2}-(L-\mu)\big{[}\hat{f}(y)-\hat{f}(x)\big{]}-\langle\mu\nabla\hat{f}(y)-L\nabla\hat{f}(x),y-x\rangle$

### v2

@@ B.2 Obtaining Mathematical Proofs
If we have a primal-dual optimal solution $(\bar{G}_{t},\bar{F}_{t},\bar{\Lambda}_{t},(\bar{\lambda}_{t,*}^{(i)}),(\bar{\lambda}_{*,t}^{(i)}))$, it holds that, for all $G_{t}\in\mathcal{S}^{2m}$ and $F_{t}\in\mathbb{R}^{2m}$,
$L(\bar{G}_{t},\bar{F}_{t},\bar{\Lambda}_{t},(\bar{\lambda}_{t,*}^{(i)}),(\bar{\lambda}_{*,t}^{(i)}))\leq L(G_{t},F_{t},\bar{\Lambda}_{t},(\bar{\lambda}_{t,*}^{(i)}),(\bar{\lambda}_{*,t}^{(i)})).$
@@
$L(\bar{G}_{t},\bar{F}_{t},\bar{\Lambda}_{t},(\bar{\lambda}_{t,*}^{(i)}),(\bar{\lambda}_{*,t}^{(i)}))=-\operatorname{Tr}(\Delta_{t}\bar{G}_{t})-\bar{F}_{t}\tilde{\Delta}=B_{t}.$
By strong duality (see Section B.3), we know that the existence of a primal-dual solution implies that $B_{t}=0$. As such, it holds that, for all $G_{t}\in\mathcal{S}^{2m}$ and $F_{t}\in\mathbb{R}^{2m}$,
$0\leq L(G_{t},F_{t},\bar{\Lambda}_{t},(\bar{\lambda}_{t,*}^{(i)}),(\bar{\lambda}_{*,t}^{(i)})),$
@@
$\operatorname{Tr}(\Delta_{t}G_{t})+F_{t}\tilde{\Delta}+\operatorname{Tr}(\bar{\Lambda}_{t}G_{t})$ $\leq\sum_{i=1}^{m}\bar{\lambda}_{t,*}^{(i)}\left(F_{t}\bm{f}_{m+i}-F_{t}\bm{f}_{i}+\operatorname{Tr}(A_{t,*}^{(i)}G_{t})\right)+\sum_{i=1}^{m}\bar{\lambda}_{*,t}^{(i)}\left(F_{t}\bm{f}_{i}-F_{t}\bm{f}_{m+i}+\operatorname{Tr}(A_{*,t}^{(i)}G_{t})\right).$ (11)
As $(\bar{G}_{t},\bar{F}_{t},\bar{\Lambda}_{t},(\bar{\lambda}_{t,*}^{(i)}),(\bar{\lambda}_{*,t}^{(i)}))$ is primal-dual optimal, in particular it is primal-dual feasible, and hence $\bar{\Lambda}_{t}\succeq 0$ and $\bar{\lambda}_{t,*}^{(i)},\bar{\lambda}_{*,t}^{(i)}\geq 0$ for all $i=1,\ldots,m$. In particular, if $G_{t}$ and $F_{t}$ are generated by functions $f_{i}\in\mathcal{F}_{\mu,L}$ through Equations (10), it holds that the right-hand side of Equation (11) is nonpositive by Theorem B.1. By recalling that $\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}=\operatorname{Tr}(\Delta_{t}G_{t})+F_{t}\tilde{\Delta}$, we write
$\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}\leq-\operatorname{Tr}(\bar{\Lambda}_{t}G_{t})\leq 0,$
where the latter follows as $\bar{\Lambda}_{t},G_{t}\succeq 0$. Specifically, one can prove that $\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}\leq 0$ by rewriting it in the form of (11), using the obtained dual optimal solution, and writing $\operatorname{Tr}(\bar{\Lambda}_{t}G_{t})$ as an appropriate sum of squares based on the Cholesky factorization of $\bar{\Lambda}_{t}$.
B.3 Strong Duality
Theorem B.2
@@ B.3 Strong Duality
$+\left(-\tilde{\Delta}+\sum_{i=1}^{m}\lambda_{*,t}^{i}(\bm{f}_{m+i}-\bm{f}_{i})+\sum_{i=1}^{m}\lambda_{*,t}^{i}(\bm{f}_{i}-\bm{f}_{m+i})\right)F_{t},$
from which we readily recover Problem (Dual) as the dual. In order to show that strong duality holds, we employ a standard Slater argument (boyd_convex_2023, ) and construct a feasible point $(G_{t},F_{t})$ to Problem (Primal) such that $G_{t}\succ 0$.
To do so, we follow the same ideas as in (taylor_smooth_2017, , Theorem 6). Specifically, leveraging the discussion in Section B.1, we shall construct $m$ functions $f_{i}\in\mathcal{F}_{\mu,L}$ and points $x_{t},x_{*}\in\mathbb{R}^{d}$ such that the matrix $P_{t}$ given by Equation (10a) is upper triangular with positive entries on its diagonal. As such, assuming $d\geq 2m$, it will hold that $G_{t}=P_{t}^{T}P_{t}\succ 0$. We assume without loss of generality that $x_{*}=0$, and set $d=2m$.
@@
where $Q_{i}^{1}\in\mathbb{R}^{(m-1)\times(m-1)}$ is a diagonal matrix with entries $2$ everywhere expect an entry of $5$ on the $i$th component, and $Q\in\mathbb{R}^{(m+1)\times(m+1)}$ is a tridiagonal matrix whose diagonal elements are $4$ and whose off diagonal elements are 1-1. We note that , $\sigma(Q_{i})\subset(1,6)$ for all $i=1,\ldots,m-1$, by the Greshgorian Theorem (quarteroni_numerical_2007, ).
We define $x_{t}=x=[y,z]$ with $y\in\mathbb{R}^{m-1}$ and $z\in\mathbb{R}^{m+1}$. Then, for $i=1,\ldots,m-1$,
$\nabla f_{i}(x_{t})=g_{t}^{(i)}=2\sum_{j=1,j\neq i}^{m-1}\bm{e}_{j}x_{j}+5\bm{e}_{i}x_{i}+Qz-\bm{e}_{m+i}x_{m+i}+b_{i}.$
@@
$b_{i}=-2\sum_{j=1,j\neq i}^{m-1}\bm{e}_{j}x_{j}+\bm{e}_{m+i}x_{m+i}-\bm{e}_{m},$
and we impose that $x>0$ and that $Qz=[1,0,\ldots,0]\in\mathbb{R}^{m+1}$ (may be shown possible by a simple induction argument). This thus gives us that $g_{t}^{(i)}=5\bm{e}_{i}x_{i}$. Moreover, as $x_{*}=0$, we have that $g_{*}^{(i)}=b_{i}$. Note that $g_{t}^{(i)}$ has a positive entry in its $i$th component and 0s everywhere else. Moreover, $g_{*}^{(i)}$ has a positive component in its $m+i$th component and 0 in all subsequent components. The vector $x_{t}-x_{*}$ also has a positive entry in component $2m$. In order to show that $P_{t}$ is upper triangular with positive diagonal entries, we thus only need to construct $g_{t}^{(m)}$.
@@
$\tilde{f}_{i}(x)=\frac{L-\mu}{5}\left(f_{i}(x)-\frac{1}{2}\|x\|^{2}\right)+\frac{\mu}{2}\|x\|^{2}.$
@@
As $f_{i}\in\mathcal{F}_{1,6}$, we have $\tilde{f}_{i}\in\mathcal{F}_{\mu,L}$. Moreover, this transformation preserves the property on $P_{t}$, hence constructing a Slater point for arbitrary $(\mu,L)$.
Appendix C SGD with Non-Uniform Sampling and Mini-Batching
@@
All the main results in our paper can be applied to variants of (SGD), such as mini-batching or nonuniform sampling. This is because these variants can be seen as instances of (SGD) applied to different but equivalent problems. Our results can be applied as a black-box, and the results follow by computing the main constants of the equivalent problem, namely $L,\mu,\sigma_{*}^{2}$. This strategy was described in gower_sgd_2019 . For the sake of completeness, we include some details for the two aforementioned instances below.
C.1 Non-Uniform Sampling
@@
where the expectation is taken with respect $\mathcal{P}$, the distribution whose density is given by the probabilities $p_{i}$. Note that the function $f$ remains unchanged, meaning that the set of minimizers is the same.
We note that Assumption 1 holds for $\mu=\min_{i}\tfrac{\mu_{i}}{np_{i}}$ and $L=\max_{i}\tfrac{L_{i}}{np_{i}}$. Moreover, one can compute that
$\hat{\sigma}_{*}^{2}=\mathbb{E}\mathopen{}\left[{\|\nabla\hat{f}_{i}(x_{*})\|^{2}}\right]\mathclose{}=\mathbb{E}\mathopen{}\left[{\frac{1}{n^{2}p_{i}^{2}}\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}\leq\frac{\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}}{n^{2}\min_{i}p_{i}^{2}}=\frac{\sigma_{*}^{2}}{n^{2}\min_{i}p_{i}^{2}},$
which is finite if Assumption 2 holds for the original problem. If we now impose that $\gamma L\in(0,2)$, we can apply all our results and obtain bounds on the iterates of (SGDp) depending on the above constants.
C.2 Mini-Batching
@@
We note that Assumption 1 holds for $\mu=\min_{i}\mu_{i}$ and $L=\max_{i}L_{i}$. Moreover, one can compute that
@@
$\sigma_{*}^{2}$ $=\mathbb{E}_{{\mathcal{B}}}\mathopen{}\left[{\|\nabla f_{B}(x_{*})\|^{2}}\right]\mathclose{}=\frac{1}{|\mathcal{B}|}\sum_{B\in\mathcal{B}}\left\|\frac{1}{b}\sum_{i\in B}\nabla f_{i}(x_{*})\right\|^{2}=\frac{1}{\binom{n}{b}\cdot b^{2}}\sum_{B\in\mathcal{B}}\sum_{i,j\in B}\langle\nabla f_{i}(x_{*}),\nabla f_{j}(x_{*})\rangle$